OpenStudy (anonymous):

factor completly 3z^2-192

OpenStudy (anonymous):

3(z^2-64)=3(z^2-8^2)=3(z-8)(z+8)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw - click "Good Answer" :)